# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ III-B3 Automated Jailbreak Prompt Generation
Liu et al. [39] present AutoDAN, a novel jailbreak attack designed to circumvent the defenses of aligned Large Language Models (LLMs), which are typically fine-tuned using human feedback to ensure safe and ethical decision-making. Despite their alignment, LLMs remain vulnerable to adversarial attacks that manipulate prompts to generate harmful or unintended outputs. Existing jailbreak techniques often face scalability issues, requiring manual prompt crafting or stealthiness problems, as they rely on token-based methods that produce semantically meaningless inputs, which can be detected through basic perplexity testing. To address these limitations, the authors introduce AutoDAN, a system that utilizes a hierarchical genetic algorithm to generate stealthy, semantically meaningful jailbreak prompts automatically. Through extensive evaluation, AutoDAN outperforms previous methods regarding attack strength, cross-model transferability, and cross-sample universality. Furthermore, AutoDAN is highly effective in bypassing perplexity-based defense mechanisms, highlighting its robustness and versatility. The findings highlight the potential for malicious actors to exploit vulnerabilities in LLMs through the automated and stealthy manipulation of prompts. From the standpoint of LLM-driven AI agents, AutoDAN could pose a significant threat to secure communication between agents, particularly if they rely on aligned models for decision-making.
@@
Yu et al. [18] present GPTFuzz, a black-box jailbreak fuzzing framework designed to automate the generation of jailbreak templates to test the safety of large language models (LLM). Although LLMs have seen widespread adoption, from casual conversations to AI-driven programming, they remain vulnerable to adversarial attacks that exploit their output to generate harmful content. Traditional jailbreak techniques rely heavily on manually crafted templates, which limits large-scale testing. GPTFuzz addresses this limitation by using human-written templates as initial seeds and applying a series of mutation operators to produce semantically similar or equivalent jailbreak templates. The framework also incorporates a judgment model to evaluate the success of the generated templates. Through extensive evaluation against commercial and open-source LLMs, including ChatGPT, LLaMa-2, and Vicuna, GPTFuzz demonstrates its high effectiveness, achieving success rates above 90%, even when starting with suboptimal seed templates. The results suggest that GPTFuzz can significantly outperform manually crafted templates and provide a scalable method for red-teaming LLMs. This attack framework can be leveraged by LLM-based AI agents to assess and secure communication channels between agents, ensuring that adversarial inputs do not compromise the integrity of multi-agent interactions. Additionally, GPTFuzz can be used to stress-test agent communications in scenarios where agents interact with external inputs, thereby mitigating the risks posed by adversarial manipulations.
@@
Fig. 6 illustrates an iterative fuzz-testing approach adapted by GPTFuzz [18] for automatically generating jailbreak prompts designed to assess and challenge the robustness of large language models (LLMs). Initially, jailbreak prompts are gathered from online sources, forming a seed pool. Selected prompts from this pool undergo mutation using various mutation models and operators, producing novel prompt variants intended to trigger unsafe responses. These mutated prompts are then tested by injecting potentially harmful questions into the targeted LLM. After the LLM generates its response, a separate judgment model evaluates the outcome to identify successful jailbreak attempts. Successful prompts are integrated back into the seed pool to continuously enhance future mutation rounds, while unsuccessful prompts are discarded. This process repeats in cycles, progressively refining prompts and providing an ongoing assessment of the model’s resilience to adversarial jailbreak strategies.
@@
Schwartz et al. [19] introduce a modular pipeline designed to automate the generation of stealthy jailbreak prompts derived from high-level content policies, thereby enhancing the content moderation capabilities of large language models (LLMs). The authors first address query inefficiency and jailbreak strength by developing the Graph of Attacks with Pruning (GAP), which utilizes strategies from previous jailbreaks to reduce the number of queries needed for a successful attack. GAP achieves a 92% success rate on GPT-3.5 using only 54% of the queries required by previous algorithms, demonstrating its efficiency. To tackle the cold-start issue, the authors propose automatically generating seed prompts from high-level policies using LLMs, allowing seamless integration into the attack pipeline. Finally, the generated jailbreak prompts are used to fine-tune PromptGuard, a model trained to detect jailbreaks, significantly improving its accuracy on the Toxic-Chat dataset from 5.1% to 93.89%. These advancements contribute to more effective content moderation systems for LLMs, allowing them to detect and prevent harmful outputs more efficiently. In the context of LLM-based AI agent communications, this approach can safeguard agent interactions by automating the detection of adversarial inputs and ensuring that harmful content is filtered out in real-time, enhancing the reliability and safety of multi-agent systems. Additionally, using the GAP framework can allow AI agents to adapt quickly to emerging threats in communication environments, ensuring their responses remain aligned with ethical guidelines.
@@
III-B4 Jailbreak‐fuzzing attack
Dong et al. [40] propose JailFuzzer, a cutting-edge fuzzing framework powered by LLM agents, which automatically crafts semantically rich, natural-language jailbreak prompts against Text-to-Image models operating in a purely black-box setting. Although T2I generators have transformed content creation by rendering detailed textual descriptions into photorealistic images, their built-in safeguards can be subverted by cleverly engineered inputs that elude detection. Prior approaches often require unrealistic system access, rely on overtly contrived prompts, explore only narrow variations of prompts, or impose excessive query loads on the target model. JailFuzzer overcomes these drawbacks by integrating three core elements: (1) a seed pool containing both benign and jailbreak prompts to bootstrap exploration; (2) a guided mutation engine that leverages LLM reasoning to produce diverse, semantically coherent prompt variants; and (3) an oracle module—also implemented via LLM agents to assess whether a given prompt successfully circumvents defenses. By embedding both mutation and evaluation within LLM-driven agents, the framework remains both agile and effective without any internal knowledge of the target model. Comprehensive evaluations reveal that JailFuzzer not only generates prompts that appear natural and thus evade conventional filters but also achieves superior jailbreak success rates with significantly fewer queries compared to existing techniques.
@@ III-B4 Jailbreak‐fuzzing attack
III-C Adversarial Example & Evasion Attacks
III-C1 An adversarial in-context learning method
Wang et al. [41] investigate the security vulnerabilities of in-context learning (ICL) in large language models (LLMs), focusing on the potential for adversarial attacks that manipulate only the demonstration examples, without altering the input itself. In-context learning enhances the performance of LLMs by providing precondition prompts in the form of data-label pairs, but it also introduces a new security concern: attackers can mislead the model by manipulating these demonstrations. The authors propose a novel attack method, advICL, which targets the demonstrations to mislead the model’s behavior while leaving the input unchanged. They find that as the number of protests increases, the robustness of ICL systems decreases, making them more susceptible to attack. Furthermore, the paper identifies an intrinsic property of demonstrations: they can be prepended to different inputs, which opens the door for a more practical and insidious attack model. To address this, the authors introduce Transferable-advICL, a version of advICL that can generate adversarial demonstrations capable of attacking unseen test inputs. Experimental results demonstrate that adversarial demonstrations generated by Transferable-advICL successfully mislead the model on inputs not included in the attack training, highlighting the broad applicability of the attack.
@@ III-C1 An adversarial in-context learning method
III-C2 Query-free adversarial attack
@@
Zhuang et al. [42] addresses the adversarial vulnerability of Text-to-Image (T2I) models, particularly focusing on Stable Diffusion, by exploring the concept of ”query-free attack generation.” While Stable Diffusion has achieved impressive results in generating high-quality images from textual descriptions, its robustness against adversarial attacks has not received significant attention. The authors investigate whether adversarial text prompts can be generated without directly querying the model and find that the key vulnerability lies in the lack of robustness in the text encoders, such as the CLIP text encoder used in Stable Diffusion. Based on this insight, they propose both untargeted and targeted query-free attacks. The untargeted attack focuses on perturbing the most influential dimensions in the text embedding space, which they term as ”steerable key dimensions.” Their experiments demonstrate that a minimal perturbation of just five characters in a text prompt can lead to a significant shift in the content of the generated image. Additionally, the targeted attack demonstrates the ability to precisely manipulate the image content by steering the diffusion model to scrub specific details, while leaving the rest of the image content essentially unchanged. These findings highlight the vulnerability of T2I models to adversarial manipulation through subtle modifications to text prompts.
@@
In deployments of AI agents utilizing LLM infrastructures, this type of query-free adversarial attack could be used to subtly alter the responses generated by agents that rely on text prompts, leading them to produce misleading or harmful outputs. By manipulating the input prompt, adversaries could steer the AI agents’ responses to focus on specific, potentially harmful content, without triggering direct model queries, making the attack more challenging to detect. This type of attack highlights the necessity for robust defenses against adversarial prompt manipulations in multi-agent systems, where agents may rely on shared prompts or external text-based instructions to make decisions.
III-C3 Multimodal adversarial attack
@@
Yang et al. [31] introduce MMA-Diffusion, a framework that poses a significant security threat to Text-to-Image (T2I) models by circumventing existing defenses against the generation of inappropriate or Not-Safe-For-Work (NSFW) content. While T2I models have seen remarkable advancements and widespread adoption, they have also inadvertently opened the door for misuse, particularly in generating harmful or inappropriate content. MMA-Diffusion stands out from previous approaches by leveraging both textual and visual modalities to bypass standard safeguards, such as prompt filters and post-hoc safety checkers, which are typically employed to prevent the generation of NSFW content. The framework highlights vulnerabilities in the current defense mechanisms of both open-source models and commercial online services, demonstrating that these defenses are insufficient in mitigating attacks that combine both textual and visual manipulation.
@@
For AI agents equipped with large language model capabilities, MMA-Diffusion can be utilized to manipulate the content generated by agents that rely on multimodal input, including text and images. By exploiting vulnerabilities in T2I models, adversaries can guide agents to produce harmful or inappropriate content without triggering safety mechanisms, particularly in systems that involve textual prompts and visual outputs. This poses a significant risk in environments where AI agents are required to generate content based on external inputs, making it essential to develop more secure and comprehensive defense strategies to safeguard against such multimodal attacks.
III-C4 Active Environment Injection Attack
@@
Chen et al. [20] introduce the Active Environment Injection Attack (AEIA), a novel threat to AI agents that arises from their inability to detect ”impostors” within their operational environment. As AI agents become more integral to task execution in mobile operating systems, their security is often overlooked, particularly in detecting malicious manipulations disguised as environmental elements. The authors highlight two critical vulnerabilities that expose AI agents to AEIA: (1) adversarial content injection in multimodal interaction interfaces, where attackers embed adversarial instructions within environmental elements, misleading the agent’s decision-making, and (2) reasoning gap vulnerabilities in the agent’s task execution process, which amplify susceptibility to AEIA attacks during reasoning. To evaluate the impact of these vulnerabilities, the authors propose AEIA-MN, an attack scheme that targets interaction vulnerabilities in mobile operating systems, specifically focusing on MLLM-based agents. Experimental results reveal that even advanced MLLMs are highly vulnerable to AEIA, with a maximum attack success rate of 93% on the AndroidWorld benchmark when both vulnerabilities are combined.
@@
Within the framework of AI agents powered by large language models, AEIA could be used to manipulate agents’ decision-making by injecting misleading environmental cues or instructions that alter the agents’ intended actions. By exploiting the agent’s vulnerability in recognizing adversarial inputs within its environment, attackers could subtly influence agent interactions, leading to incorrect or harmful outputs in multi-agent systems.
III-C5 Context manipulation attack
@@
Patlan et al. [43] explore the security vulnerabilities of AI agents integrated within Web3 ecosystems, specifically in blockchain-based financial systems, where these agents interact with decentralized protocols and immutable smart contracts. While AI agents offer significant autonomy and openness in such environments, they expose new, underexplored security risks. The authors introduce context manipulation as a comprehensive attack vector, exploiting unprotected surfaces within the agent’s input channels, memory modules, and external data feeds. Through empirical analysis of ElizaOS, a decentralized AI agent framework for automated Web3 operations, the paper demonstrates how adversaries can manipulate context by injecting malicious instructions into prompts or historical interaction records. This manipulation can lead to unintended asset transfers or protocol violations, with potentially devastating financial consequences. To evaluate the severity of these vulnerabilities, the authors propose CrAIBench, a Web3 domain-specific benchmark designed to assess the robustness of AI agents against context manipulation attacks across more than 150 realistic blockchain tasks, including token transfers, trading, and cross-chain interactions, as well as over 500 attack test cases. The results highlight that while prompt-based defenses are ineffective against context manipulation, fine-tuning-based defenses provide more robust protection by substantially reducing attack success rates while maintaining utility in single-step tasks. These findings underscore the need for secure and fiduciary-responsible AI agents that can mitigate adversarial threats in Web3 ecosystems.
@@
Yang et al.[44] 2024 Backdoor Agent Backdoor Attack (BadAgent) LLM‐Based Intelligent Agents Calls for secure fine‐tuning protocols; no specific mitigation tested Agents remain vulnerable even when fine‐tuned on trustworthy data; backdoor triggers manipulate intermediate reasoning
@@
Cai et al.[45] 2022 Backdoor Prompt‐Level Backdoors (BadPrompt) Continuous prompt‐based learning models Two‐step trigger generation and optimization (TCG & ATO) BadPrompt generates subtle backdoor triggers in few‐shot learning; minimal semantic similarity to non‐target samples
@@
Yao et al.[46] 2024 Backdoor Prompt‐Based Backdoors (PoisonPrompt) Hard & Soft Prompt Methods on LLMs Demonstrates need for prompt validation in PraaS pipelines PoisonPrompt compromises prompt integrity across three methods, six datasets, three LLMs with subtle triggers
@@
Wang et al.[47] 2024 Backdoor Model‐Parameter Backdoors (BadAgent) LLM Agents with external tools Highlights requirement for robust vetting of third‐party models Backdoor remains effective after fine‐tuning; agents misused to perform harmful operations with simple triggers
@@
Huang et al.[21] 2023 Backdoor Composite Backdoor Attack (CBA) LLaMA‐7B; multimodal tasks Introduces multi‐trigger key embedding; negative sample strategy CBA achieves 100% ASR with 3% poisoning on Emotion dataset; false trigger rate <<2.06%, negligible clean accuracy drop
@@
Zhu et al.[22] 2025 Backdoor Encrypted Multi‐Backdoor Implantation (DemonAgent) LLM‐Based Agents Dynamic encryption & fragment merging; AgentBackdoorEval dataset Nearly 100% ASR with 0% detection; dynamic encryption hides backdoor signals in benign content
@@
Alber et al.[48] 2025 Data‐Poisoning Medical Misinformation Poisoning The Pile (medical subset) Biomedical knowledge‐graph screening Only 0.001% token poisoning raises harmful output likelihood; knowledge‐graph screening captures 91.9% harmful content (F1 = 85.7%)
@@
Zou et al.[23] 2024 Data‐Poisoning Retrieval Poisoning (PoisonedRAG) RAG Systems (millions of texts) Evaluates and finds existing defenses inadequate 90% ASR by injecting five malicious texts; black‐box/white‐box optimization formulations
@@
Zhang et al.[49] 2024 Data‐Poisoning Retrieval Poisoning RAG‐Augmented Applications Position bias elimination strategies 88.33% success in controlled experiments; 66.67% success in real‐world applications via visually indistinguishable malicious documents
@@
Wallace et al.[50] 2020 Data‐Poisoning Gradient‐Based Backdoor Poisoning Email Spam Classifier; various NLP tasks Three potential defenses (trade‐offs in accuracy or require extra annotation) 50 poison examples suffice for a sentiment model to misclassify any input with “James Bond”; stealthy gradient‐based crafting
@@
Fang et al.[51] 2020 Data‐Poisoning Federated Local Model Poisoning Federated Learning Methods (FedAVG, FedOPT, FedNOVA) Tests four Byzantine‐robust FL methods; finds existing defenses insufficient Attacks degrade performance of Byzantine‐robust methods; one defense works in limited cases
@@
Dong et al.[52] 2025 Data‐Poisoning Memory‐Poisoning Attack (MINJA) LLM Agent Memory Banks Progressive shortening strategy; no direct defense proposed MINJA injects malicious records via queries to cause harmful reasoning; robust across various agents with minimal interaction
@@
Model Compromise Attacks Backdoor Attacks BadPrompt (Cai et al. [45]) PoisonPrompt (Yao et al. [46]) BadAgent (Wang et al. [47]) DemonAgent (Zhu et al. [22]) Composite Backdoor (Huang et al. [21]) Data- Poisoning Medical Misinformation (Alber et al. [48]) PoisonedRAG (Zou et al. [23]) Retrieval Poisoning (Zhang et al. [49]) Gradient-Based Backdoor Poisoning (Wallace et al. [50]) FL Local Model Poisoning (Fang et al. [51]) Memory- Poisoning MINJA (Dong et al. [52])

### v2

@@ III-C3 Automated Jailbreak Prompt Generation
Liu et al. [liu2023autodan] present AutoDAN, a novel jailbreak attack designed to circumvent the defenses of aligned Large Language Models (LLMs), which are typically fine-tuned using human feedback to ensure safe and ethical decision-making. Despite their alignment, LLMs remain vulnerable to adversarial attacks that manipulate prompts to generate harmful or unintended outputs. Existing jailbreak techniques often face scalability issues, requiring manual prompt crafting or stealthiness measures, as they rely on token-based methods that produce semantically meaningless inputs that can be detected through basic perplexity testing. To address these limitations, the authors introduce AutoDAN, a system that uses a hierarchical genetic algorithm to automatically generate stealthy, semantically meaningful jailbreak prompts. Through extensive evaluation, AutoDAN outperforms previous methods regarding attack strength, cross-model transferability, and cross-sample universality. Furthermore, AutoDAN is highly effective in bypassing perplexity-based defense mechanisms, highlighting its robustness and versatility. The findings highlight the potential for malicious actors to exploit LLM vulnerabilities through automated, stealthy prompt manipulation. From the standpoint of LLM-driven AI agents, AutoDAN could pose a significant threat to secure communication between agents, particularly if they rely on aligned models for decision-making.
@@
Yu et al. [yu2023gptfuzzer] present GPTFuzz, a black-box jailbreak fuzzing framework designed to automate the generation of jailbreak templates to test the safety of large language models (LLM). Although LLMs have seen widespread adoption—from casual conversations to AI-driven programming—they remain vulnerable to adversarial attacks that exploit their outputs to generate harmful content. Traditional jailbreak techniques rely heavily on manually crafted templates, which limit large-scale testing. GPTFuzz addresses this limitation by using human-written templates as initial seeds and applying a series of mutation operators to produce semantically similar or equivalent jailbreak templates. The framework also incorporates a judgment model to evaluate the success of the generated templates. Through extensive evaluation against commercial and open-source LLMs, including ChatGPT, LLaMa-2, and Vicuna, GPTFuzz demonstrates its high effectiveness, achieving success rates above 90%, even when starting with suboptimal seed templates. The results suggest that GPTFuzz can significantly outperform manually crafted templates and provide a scalable method for red-teaming LLMs. This attack framework can be leveraged by LLM-based AI agents to assess and secure communication channels, ensuring that adversarial inputs do not compromise the integrity of multi-agent interactions. Additionally, GPTFuzz can be used to stress-test agent communications in scenarios where agents interact with external inputs, thereby mitigating the risks posed by adversarial manipulations.
@@
Fig. 6 illustrates an iterative fuzz-testing approach adapted by GPTFuzz [yu2023gptfuzzer] for automatically generating jailbreak prompts designed to assess and challenge the robustness of large language models (LLMs). Initially, jailbreak prompts are gathered from online sources to form a seed pool. Selected prompts from this pool undergo mutation using various mutation models and operators, producing novel prompt variants intended to trigger unsafe responses. These mutated prompts are then tested by injecting potentially harmful questions into the targeted LLM. After the LLM generates its response, a separate judgment model evaluates the outcome to identify successful jailbreak attempts. Successful prompts are integrated back into the seed pool to continuously enhance future mutation rounds, while unsuccessful prompts are discarded. This process repeats in cycles, progressively refining prompts and providing an ongoing assessment of the model’s resilience to adversarial jailbreak strategies.
@@
Schwartz et al. [schwartz2025graph] introduce a modular pipeline designed to automate the generation of stealthy jailbreak prompts derived from high-level content policies, thereby enhancing the content moderation capabilities of large language models (LLMs). The authors first address query inefficiency and jailbreak strength by developing the Graph of Attacks with Pruning (GAP), which leverages strategies from prior jailbreaks to reduce the number of queries required for a successful attack. GAP achieves a 92% success rate on GPT-3.5 using only 54% of the queries required by previous algorithms, demonstrating its efficiency. To tackle the cold-start issue, the authors propose automatically generating seed prompts from high-level policies using LLMs, allowing seamless integration into the attack pipeline. Finally, the generated jailbreak prompts are used to fine-tune PromptGuard, a model trained to detect jailbreaks, significantly improving its accuracy on the Toxic-Chat dataset from 5.1% to 93.89%. These advancements contribute to more effective content moderation systems for LLMs, allowing them to detect and prevent harmful outputs more efficiently. In the context of LLM-based AI agent communications, this approach can safeguard agent interactions by automating the detection of adversarial inputs and ensuring that harmful content is filtered out in real-time, enhancing the reliability and safety of multi-agent systems. Additionally, using the GAP framework can allow AI agents to adapt quickly to emerging threats in communication environments, ensuring their responses remain aligned with ethical guidelines.
@@
III-C4 Jailbreak‐fuzzing attack
Dong et al. [dong2025fuzz] propose JailFuzzer, a cutting-edge fuzzing framework powered by LLM agents, which automatically crafts semantically rich, natural-language jailbreak prompts against Text-to-Image models operating in a purely black-box setting. Although T2I generators have transformed content creation by rendering detailed textual descriptions into photorealistic images, their built-in safeguards can be subverted by carefully crafted inputs that evade detection. Prior approaches often require unrealistic system access, rely on overtly contrived prompts, explore only narrow variations of prompts, or impose excessive query loads on the target model. JailFuzzer overcomes these drawbacks by integrating three core elements: (1) a seed pool containing both benign and jailbreak prompts to bootstrap exploration; (2) a guided mutation engine that leverages LLM reasoning to produce diverse, semantically coherent prompt variants; and (3) an oracle module—also implemented via LLM agents to assess whether a given prompt successfully circumvents defenses. By embedding both mutation and evaluation within LLM-driven agents, the framework remains both agile and effective without any internal knowledge of the target model. Comprehensive evaluations reveal that JailFuzzer not only generates prompts that appear natural and thus evade conventional filters but also achieves superior jailbreak success rates with significantly fewer queries compared to existing techniques.
@@ III-C4 Jailbreak‐fuzzing attack
III-D Adversarial Example & Evasion Attacks
III-D1 An adversarial in-context learning method
Wang et al. [wang2023adversarial] investigate the security vulnerabilities of in-context learning (ICL) in large language models (LLMs), focusing on the potential for adversarial attacks that manipulate only the demonstrating the input itself. In-context learning enhances LLM performance by providing precondition prompts in the form of data-label pairs, but it also introduces a new security concern: attackers can mislead the model by manipulating these demonstrations. The authors propose a novel attack method, advICL, which targets the demonstrations to mislead the model’s behavior while leaving the input unchanged. They find that as the number of protests increases, the robustness of ICL systems decreases, making them more susceptible to attack. Furthermore, the paper identifies an invariant of advICL: it can be prepended to different inputs, which opens the door to a more practical and insidious attack model. To address this, the authors introduce Transferable-advICL, a version of advICL that can generate adversarial demonstrations capable of attacking unsattack’s broad applicability. Results demonstrate that adversarial demonstrations generated by Transferable-advICL successfully mislead the model on inputs not included in the attack training, highlighting the attack’s broad applicability.
@@ III-D1 An adversarial in-context learning method
III-D2 Query-free adversarial attack
@@
Zhuang et al. [zhuang2023pilot] address the adversarial vulnerability of Text-to-Image (T2I) models, particularly focusing on Stable Diffusion, by exploring the concept of ”query-free attack generation.” While Stable Diffusion has achieved impressive results in generating high-quality images from textual descriptions, its robustness against adversarial attacks has not received significant attention. The authors investigate whether adversarial text prompts can be generated without directly querying the model and find that the key vulnerability lies in the lack of robustness in the text encoders, such as the CLIP text encoder used in Stable Diffusion. Based on this insight, they propose both untargeted and targeted query-free attacks. The untargeted a5 characters in a text prompt can lead to a significant shift in the generated image’s contentrm as ”steerable key dimensions.” Their experiments demonstrate that a minimal perturbation of just 5 characters in a text prompt can lead to a significant shift in the generated image’s content. Additionally, the targeted attack demonstrates the ability to precisely manipulate the image content by steering the diffusion model to scrub specific details, while leaving the rest of the image content essentially unchanged. These findings highlight the vulnerability of T2I models to adversarial manipulation through subtle modifications to text prompts.
@@
In deployments of AI agents utilizing LLM infrastructures, this type of query-free adversarial attack could be used to subtly alter the responses generated by agents that rely on text prompts, leading them to produce misleading or harmful outputs. By manipulating the input prompt, adversaries underscore the need for robust defenses against adversarial prompt manipulation that do not trigger direct model queries, thereby making the attack more challenging to detect. This type of attack underscores the need for robust defenses against adversarial prompt manipulation in multi-agent systems, where agents may rely on shared prompts or external text-based instructions to make decisions.
III-D3 Multimodal adversarial attack
@@
Yang et al. [yang2024mma] introduce MMA-Diffusion, a framework that poses a significant security threat to Text-to-Image (T2I) models by circumventing existing defenses against the generation of inappropriate or Not-Safe-For-Work (NSFW) content. While T2I models have seen remarkable advancements and widespread adoption, they have also inadvertently opened the door for misuse, particularly in generating harmful or inappropriate content. MMA-Diffusion stands out from previous approaches by leveraging both textual and visual modalities to bypass standard safeguards, such as prompt filters and post-hoc safety checkers, which are typically employed to prevent the generation of NSFW content. The framework highlights vulnerabilities in the current defenses of both open-source models and commercial online services, demonstrating that these defenses are insufficient to mitigate attacks that combine textual and visual manipulation.
@@
For AI agents equipped with large language model capabilities, MMA-Diffusion can be utilized to manipulate the content generated by agents that rely on recognizing adversarial inputs in their environment. Attackers can use T2I models, and adversaries can guide agents to produce harmful or inappropriate content without triggering safety mechanisms, particularly in systems that involve textual prompts and visual outputs. This poses a significant risk in environments where AI agents generate content based on external inputs, making it essential to develop more secure and comprehensive defense strategies to safeguard against such multimodal attacks.
III-D4 Active Environment Injection Attack
@@
Chen et al. [chen2025aeia] introduce the Active Environment Injection Attack (AEIA), a novel threat that exploits their inability to detect ”impostors” within their operational environment. As AI agents become more integral to task execution in mobile operating systems, their security is often overlooked, particularly in detecting malicious manipulations disguised as environmental elements. The authors highlight two critical vulnerabilities that expose AI agents to AEIA: (1) adversarial content injection in multimodal interaction interfaces, where attackers embed adversarial instructions within environmental elements, misleading the agent’s decision-making, and (2) reasoning gap vulnerabilities in the agent’s task execution process, which amplify susceptibility to AEIA attacks during reasoning. To evaluate the impact of these vulnerabilities, the authors propose AEIA-MN, an attack that targets interaction vulnerabilities. They show that while prompt-based defenses are ineffective against context manipulation, fine-tuning-based defenses provide more robust protection against AEIA, achieving a maximum attack success rate of 93% on the AndroidWorld benchmark when both vulnerabilities are combined.
@@
Within the framework of AI agents powered by large language models, AEIA could be used to manipulate agents’ decision-making by injecting misleading environmental cues or instructions that alter the agents’ intended actions. By exploiting the agent’s vulnerability to recognizing adversarial inputs in its environment, attackers can subtly influence agent interactions, leading to incorrect or harmful outputs in multi-agent systems.
III-D5 Context manipulation attack
@@
Patlan et al. [patlan2025real] explore the security vulnerabilities of AI agents integrated within Web3 ecosystems, specifically in blockchain-based financial systems, where these agents interact with decentralized protocols and immutable smart contracts. While AI agents offer significant autonomy and openness in such environments, they expose new, underexplored security risks. The authors introduce context manipulation as a comprehensive attack vector that exploits unprotected surfaces within the agent’s input channels, memory modules, and external data feeds. Through empirical analysis of ElizaOS, a decentralized AI agent framework for automated Web3 operations, the paper demonstrates how adversaries can manipulate context by injecting malicious instructions into prompts or historical interaction records. This manipulation can lead to unintended asset transfers or protocol violations, with potentially devastating financial consequences. To evaluate the severity of these vulnerabilities, the authors propose CrAIBench, a Web3 domain-specific benchmark designed to assess the robustness of AI agents against context manipulation attacks across more than 150 realistic blockchain tasks, including token transfers, trading, and cross-chain interactions, and over 500 attack test cases. The results show that while prompt-based defenses are ineffective against context manipulation, fine-tuning-based defenses provide more robust protection, substantially reducing attack success rates while maintaining utility in single-step tasks. These findings underscore the need for secure and fiduciary-responsible AI agents that can mitigate adversarial threats in Web3 ecosystems.
@@
Yang et al.[yang2024watch] 2024 Backdoor Agent Backdoor Attack (BadAgent) LLM‐Based Intelligent Agents Calls for secure fine‐tuning protocols; no specific mitigation tested Agents remain vulnerable even when fine‐tuned on trustworthy data; backdoor triggers manipulate intermediate reasoning
@@
Cai et al.[cai2022badprompt] 2022 Backdoor Prompt‐Level Backdoors (BadPrompt) Continuous prompt‐based learning models Two‐step trigger generation and optimization (TCG & ATO) BadPrompt generates subtle backdoor triggers in few‐shot learning; minimal semantic similarity to non‐target samples
@@
Yao et al.[yao2024poisonprompt] 2024 Backdoor Prompt‐Based Backdoors (PoisonPrompt) Hard & Soft Prompt Methods on LLMs Demonstrates need for prompt validation in PraaS pipelines PoisonPrompt compromises prompt integrity across three methods, six datasets, three LLMs with subtle triggers
@@
Wang et al.[wang2024badagent] 2024 Backdoor Model‐Parameter Backdoors (BadAgent) LLM Agents with external tools Highlights requirement for robust vetting of third‐party models Backdoor remains effective after fine‐tuning; agents misused to perform harmful operations with simple triggers
@@
Huang et al.[huang2023composite] 2023 Backdoor Composite Backdoor Attack (CBA) LLaMA‐7B; multimodal tasks Introduces multi‐trigger key embedding; negative sample strategy CBA achieves 100% ASR with 3% poisoning on Emotion dataset; false trigger rate <<2.06%, negligible clean accuracy drop
@@
Zhu et al.[zhu2025demonagent] 2025 Backdoor Encrypted Multi‐Backdoor Implantation (DemonAgent) LLM‐Based Agents Dynamic encryption & fragment merging; AgentBackdoorEval dataset Nearly 100% ASR with 0% detection; dynamic encryption hides backdoor signals in benign content
@@
Alber et al.[alber2025medical] 2025 Data‐Poisoning Medical Misinformation Poisoning The Pile (medical subset) Biomedical knowledge‐graph screening Only 0.001% token poisoning raises harmful output likelihood; knowledge‐graph screening captures 91.9% harmful content (F1 = 85.7%)
@@
Zou et al.[zou2024poisonedrag] 2024 Data‐Poisoning Retrieval Poisoning (PoisonedRAG) RAG Systems (millions of texts) Evaluates and finds existing defenses inadequate 90% ASR by injecting five malicious texts; black‐box/white‐box optimization formulations
@@
Zhang et al.[zhang2024human] 2024 Data‐Poisoning Retrieval Poisoning RAG‐Augmented Applications Position bias elimination strategies 88.33% success in controlled experiments; 66.67% success in real‐world applications via visually indistinguishable malicious documents
@@
Wallace et al.[wallace2020concealed] 2020 Data‐Poisoning Gradient‐Based Backdoor Poisoning Email Spam Classifier; various NLP tasks Three potential defenses (trade‐offs in accuracy or require extra annotation) 50 poison examples suffice for a sentiment model to misclassify any input with “James Bond”; stealthy gradient‐based crafting
@@
Fang et al.[fang2020local] 2020 Data‐Poisoning Federated Local Model Poisoning Federated Learning Methods (FedAVG, FedOPT, FedNOVA) Tests four Byzantine‐robust FL methods; finds existing defenses insufficient Attacks degrade performance of Byzantine‐robust methods; one defense works in limited cases
@@
Dong et al.[dong2025practical] 2025 Data‐Poisoning Memory‐Poisoning Attack (MINJA) LLM Agent Memory Banks Progressive shortening strategy; no direct defense proposed MINJA injects malicious records via queries to cause harmful reasoning; robust across various agents with minimal interaction
@@
Model Compromise Attacks Backdoor Attacks BadPrompt (Cai et al. [cai2022badprompt]) PoisonPrompt (Yao et al. [yao2024poisonprompt]) BadAgent (Wang et al. [wang2024badagent]) DemonAgent (Zhu et al. [zhu2025demonagent]) Composite Backdoor (Huang et al. [huang2023composite]) Data- Poisoning Medical Misinformation (Alber et al. [alber2025medical]) PoisonedRAG (Zou et al. [zou2024poisonedrag]) Retrieval Poisoning (Zhang et al. [zhang2024human]) Gradient-Based Backdoor Poisoning (Wallace et al. [wallace2020concealed]) FL Local Model Poisoning (Fang et al. [fang2020local]) Memory- Poisoning MINJA (Dong et al. [dong2025practical])